OpenStudy (anonymous):

21 quarters out of 186 quarters as a fractionin simplest form

OpenStudy (anonymous):

\[\frac{21}{186}=\frac{7}{62} \]

OpenStudy (anonymous):

First you should set up the fraction so \[\frac{ 21 }{186 }\]. Then to put the fraction in simplest form you have to look at the numerator and denominator and see if there is a number that can divide both evenly.

OpenStudy (anonymous):

One way to do this is to look at the factors of the numerator and see if they can evenly divide the denominator. So the factors of 21 are 1, 3, 7, 21. Then see if any of those factors divide 186. so we get: \[\frac{ 186 }{ 21 }\] = 8.85 \[\frac{ 186 }{ 7 }\] =26.57 \[\frac{ 186 }{ 3 }\]=62 \[\frac{ 186 }{ 1 }\] =186 The only factors of 21 that evenly divide into 186 are 1 and 3. Since 1 gives us the same number back it doesn't help us put the fraction in simplest form so we use 3. \[\frac{ 21}{ 3 }\]=7 and \[\frac{ 186 }{ 3 }\] =62 so our new simplified fraction is \[\frac{ 7}{ 62 }\]